OpenStudy (anonymous):

Eight less than a number doubled is twenty more than the number. What is the number?

OpenStudy (anonymous):

Let the number be X. 2(X - 8) = 20 + X X = ?

OpenStudy (anonymous):

So Helpful Thnaks